Grade 12th passAlgebra

The ages of Rahul and Haroon are in the ratio 5:7. Four years later the sum of their ages will be 56 years. What are their present ages?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the ages of Rahul and Haroon be 5x and 7x.
Four years later,
The ages of Rahul and Haroon will be (5x + 4) and (7x + 4) respectively. According to the question,
(5x+4)+(7x+4)=56
⇒5x+4+7x+4=56
⇒12x+8=56
⇒12x=56-8
⇒12x=48
⇒x=48/12
⇒x=4
Therefore, Present age of Rahul = 5x = 5×4 = 20
And, present age of Haroon = 7x = 7×4 = 28

Thanks